OpenStudy (anonymous):

Solve: 8x - 2(x + 1) = 7x + 8 Answer -10 10 6 -6 Solve: 8x - 2(x + 1) = 7x + 8 Answer -10 10 6 -6 @Mathematics

OpenStudy (anonymous):

start working this out first 8x - 2(x + 1) = 7x + 8 start with this part -2(x + 1)

OpenStudy (anonymous):

8x - 2(x + 1) = 7x + 8 8x - 2x - 2 = 7x + 8 6x - 2 = 7x + 8 -2 = x + 8 -10 = x

OpenStudy (anonymous):

8x - 2(x + 1) = 7x + 8 8x - 2x - 2 = 7x + 8 6x - 2 = 7x + 8 -2 = x + 8 -10 = x

OpenStudy (anonymous):

wht is ur best approach @Bamboo350? where did u got stuck?